OpenStudy (anonymous):

PLEASE HELP ....In the unbalanced equation below, what element is being reduced? MnO2 + HCl → H2O + MnCl2 + Cl2 A. Mn B. O C. H D. Cl

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@jhonyy9 someone plz hlp

OpenStudy (anonymous):

can someomne confirm a trying to get a perfect 100

thomaster (thomaster):

write the half reactions, then see which one gets reduced. reduction = gaining an electron

OpenStudy (anonymous):

so it would be O correct

thomaster (thomaster):

no

OpenStudy (anonymous):

ty for helping me

OpenStudy (anonymous):

im lost n have 7 min to finish

OpenStudy (anonymous):

well at the end its gains cl..

thomaster (thomaster):

okay you need to identify the oxidation state of every atom

OpenStudy (anonymous):

ok

thomaster (thomaster):

Mn and Cl on both sides of the arrow

thomaster (thomaster):

whats the oxidation state of Mn on the left?

OpenStudy (anonymous):

o2

thomaster (thomaster):

you can calculate this since O is 2-, you have 2 oxygen atoms so 4- to cancel this out, the Mn has to be 4+. Do you understand that?

OpenStudy (anonymous):

no im sorry im so lost and i have 2 min to turn it in just confirm itss A?

thomaster (thomaster):

Cl is 1- On the right side of the arrow, Mn is with 2x Cl so the Cl is 2- . the Mn has to be 2+ The Mn goes from 4+ to 2+, so it gains 2 electrons, this it the element that gets reduced So A is correct :D

OpenStudy (anonymous):

thank u!!!!!! tahts wut i said in beginning lol thank u!!

thomaster (thomaster):

yes but you have to understand why. It's useless to just tell you it's A. Do you understand the stuff i just explained?

OpenStudy (anonymous):

yea i understand some but im somewhat lost in chem especially talking it over the internet for a high school program.. i have my ged bc of my work but i need a "official diploma" for my dad to be happy

thomaster (thomaster):

okay :P Just remember O= 2- H= 1+ Cl = 1- Mn can have various oxidation states. When you see HCl, the total oxidation state has to be 0. The hydrogen is 1+ and the chlorine is -1. these 2 cancel eachother out as 1-1 = 0 Now for MnO2, you know oxygen is 2- You have 2 oxygen atoms so the total is 4- To cancel the other half out the manganese has to be 4+ :P for the MnCl2 it works the same but not its with chlorine which is 1- 2x chloring = 2- to cancel the other half out the manganese has to be 2+ When an element gains an electron, the oxidation state decreases by one because electrons have a negative charge. So when the oxidation state goes from +4 to +2 you know it gains 2 electrons. and that is reduction :D \(\overline{\underline{\LARGE{\color{gold}{\star~}}\Large\tt\color{green}{I\;Hope\;this\;Helps!}\LARGE{\color{gold}{~\star}}}}\)

OpenStudy (nincompoop):

did you figure things out?

OpenStudy (nincompoop):

think of those oxidation states like money, they get reduced if they became lower. it is easy for the (+) oxidation states, but how about when you just have negative (-)? let us say from 2- to 3-, see mathematically the value is lower for 3- so it still follows the idea that it got reduced. But before you solve anything, just write down the reaction and products first, then put their oxidation states as you balance them.

OpenStudy (anonymous):

yea thank u guys for your help!! i really appriciate all the extra help!!!!

OpenStudy (anonymous):

i always mess up writing them dwn wrong